OpenStudy (anonymous):

The temperature outside is 2° Celsius. The formula for degrees Celsius is: C = 9 over 5 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 2 degrees Celsius is.

OpenStudy (texaschic101):

Part 1 : C = 9/5(F - 32) --- multiply both sides by 5/9 5/9C = F - 32 -- add 32 to both sides 5/9C + 32 = F Part 2 : 5/9C + 32 = F 5/9C + 32 = 2 5/9C = 2 - 32 5/9C = -30 C = -30/(5/9) C = -30 * 9/5 C = - 270/5 C = - 54